A Complete Bibliography of Publications in IEEE Transactions on Privacy

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

26 August 2024 Version 1.01

Title word cross-reference

Anomaly [AMWV24]. Challenge [AMWV24]. Cover [Ano24]. Detection [AMWV24]. Editorial [Vai24]. Enhanced [AMWV24]. Federated [AMWV24]. Front [Ano24]. IEEE [Vai24]. Introduction [Vai24]. Learning [AMWV24]. PETs [AMWV24]. Privacy [AMWV24, Vai24]. Privacy-Enhanced [AMWV24]. Prize [AMWV24]. Transactions [Vai24]. U.K [AMWV24]. U.S. [AMWV24]. via [AMWV24].

References

Asif:2024:UUK

[AMWV24] Hafiz Asif, Sitao Min, Xinyue Wang, and Jaideep Vaidya. U.S.– U.K. PETs Prize Challenge: Anomaly detection via privacyenhanced federated learning.

IEEE Transactions on Privacy, 1 (??):3–18, 2024. ISSN 2836-208X (electronic).

Anonymous:2024:FC

[Ano24] Anonymous. Front cover. IEEE Transactions on Privacy, 1(??): C1, 2024. CODEN ???? ISSN 2836-208X (electronic).

Vaidya:2024:EII

[Vai24] Jaideep Vaidya. Editorial introduction to the IEEE Transactions on Privacy. IEEE Transactions on Privacy, 1(??):1–2, 2024. CO-DEN ???? ISSN 2836-208X (electronic).